# Convexified Message-Passing Graph Neural Networks

### v1

@@ G.1 Dataset Descriptions
G.2 Additional Implementation Details
@@
Hardware and Software. All experiments could fit on one GPU at a time. Most experiments were run on a server with 4 NVIDIA L40S GPUs. We ran all of our experiments in Python, using the PyTorch framework [34] (license URL). We also make use of PyTorch Geometric (PyG) [14] (MIT License) for experiments with graph data.
@@
Implementation Details. Particularly, all models were implemented using PyTorch Geometric (PyG) [14]. For GATv2 and GraphGPS, we use the official PyG implementations of their original papers, while for other baselines we use their PyG implementations, as standard in prior works (see, e.g., [32, 36]). Each neural architecture consists of $L\in\{2,6\}$ layers of the considered baseline, followed by a Set2Set global pooling [46] to aggregate the learned node-level representations to a graph-level embedding and a two-layer MLP for the final classification. The embedding dimension of each hidden layer is set to $64$, while non-convex models use a ReLU activation function between layers. For convex models, we use the Gaussian RBF kernel $\kappa(\mathbf{z},\mathbf{z}^{\prime}):=\exp(-\gamma\|\mathbf{z}-\mathbf{z}^{\prime}\|_{2}^{2})$ with $\gamma=0.2$, and compute an approximate kernel matrix via Nyström approximation [48] with dimension $P=64$ (Recall Section 4.4). As we want to evaluate the performance of GNN models in isolation, note that we do not use DropOut or batch normalization. Further, since we do not focus on the impact of attention heads on performance, all convex and non-convex versions of GAT, GATv2, GraphTrans and GraphGPS use their default configurations of attention heads.
Appendix H Additional Experimental Results
H.1 Heatmap of Classification Accuracy
Figure 2 provides a heatmap summarizing the accuracy of all models across the different datasets. It complements Table 1 by visually reinforcing that convex models generally outperform their non-convex counterparts across nearly all datasets, with consistent and major improvements even for models like GATv2 and GraphGPS. This reaffirms the strength and generality of our convexification framework. For example, on MUTAG, the convex CGIN model reaches 83.2% accuracy, outperforming the two-layer (60.0%) and six-layer (61.2%) GINs. Similarly, CGraphGPS achieves 83.7%, exceeding the best non-convex variant by over 30%. This trend holds across other datasets like PTC-MR, PROTEINS, and NCI1, where convex versions of CGraphSAGE, CGATv2, and CGraphGPS outperform non-convex baselines by 10–40%.
@@ H.1 Heatmap of Classification Accuracy
H.2 Bar Plots

### v2

@@ H.1 Dataset Descriptions
ogbg-molhiv 41,127 25 27 2
H.2 Additional Implementation Details
@@
Hardware and Software. All experiments could fit on one GPU at a time. Most experiments were run on a server with 4 NVIDIA L40S GPUs. We ran all of our experiments in Python, using the PyTorch framework (Paszke et al., 2019) (license URL). We also make use of PyTorch Geometric (PyG) (Fey and Lenssen, 2019) (MIT License) for experiments with graph data.
@@
Implementation Details. Particularly, all models were implemented using PyTorch Geometric (PyG) (Fey and Lenssen, 2019). For GATv2 and GraphGPS, we use the official PyG implementations of their original papers, while for other baselines we use their PyG implementations, as standard in prior works (see, e.g., (Morris et al., 2019; Rampášek et al., 2022)). In a high level, each neural architecture with $L\in\{2,6\}$ layers consists of an input encoder, stacked GNN layers with edge encoders, batch normalization as well as optional dropout and residual connections, followed by a global pooling step and a final output encoder (See Appendix H.3 for more details). The embedding dimension of each hidden layer is set to $32$, while non-convex models use a ReLU activation function between layers. For convex models, we use the Gaussian RBF kernel $\kappa(\mathbf{z},\mathbf{z}^{\prime}):=\exp(-\gamma\|\mathbf{z}-\mathbf{z}^{\prime}\|_{2}^{2})$ with $\gamma=0.2$, and compute an approximate kernel matrix via Nyström approximation (Williams and Seeger, 2001) with dimension $P=32$ (Recall Section 4.4 in the full paper). Further, since we do not focus on the impact of attention heads on performance, all convex and non-convex versions of GAT, GATv2, GraphTrans and GraphGPS use their default configurations of attention heads.
Optimization. All models are trained over $200$ epochs with the Adam optimizer (Kingma and Ba, 2015) using the cross entropy loss, a starting learning rate of $10^{-3}$, a minimum learning rate of $10^{-6}$ and a weight decay of $10^{-5}$. We use a ReduceLROnPlateau scheduler, which cuts the learning rate in half with a patience of $20$ epochs, and training ends when we reach the minimum learning rate.
H.3 Detailed Implementation Details of Neural Architectures
Node Input Encoder.
The first layer of each neural architecture consists of an input encoder implemented using an MLP, which transforms the raw data captured by the graph signal into a suitable representation for the GNN to process. Specifically:
  • If raw node features are available, then we an MLP whose input dimension is the number of features per node and its output dimension is the embedding dimension of $32$.
  • If node features are absent, we use a learnable DiscreteEncoder that maps integer node identifiers to vectors to the embedding dimension of $32$.
Edge Feature Encoders.
We construct a separate edge encoder for each GNN layer. Each edge encoder maps raw edge features to the hidden dimension:
  • If edge features are provided, then we use a single-layer MLP.
  • If edge features are absent, we use a DiscreteEncoder.
GNN Layers.
We use a stack of $L$ GNN layers, each operating over the hidden feature dimension. Each GNN layer $\ell$ takes node features encoded using the input encoder, edge features encoded using the edge encoder corresponding to layer $\ell$, and an edge index matrix. Batch normalization is then applied to the output of the GNN layer, followed by a ReLU activation function. An optional dropout and a residual connection are then applied. Each layer thus has the following general form:
$\text{GNN Layer}\rightarrow\text{BatchNorm}\rightarrow\text{ReLU}\rightarrow\text{Dropout (Optional)}\rightarrow\text{Residual Connection (Optional)}$
Global Pooling.
After all GNN layers, we aggregate node features to obtain graph-level representations using add pooling to aggregate the learned node-level representations to a graph-level embedding and a two-layer MLP for the final classification.
Output MLP.
The final graph representation is passed through a two-layer MLP for the final classification. The first layer applies a linear projection followed by optional normalization and a nonlinearity. The second layer outputs logits or regression scores without an activation function.
Appendix I Additional Experimental Results
I.1 Heatmap of Classification Accuracy
@@
Figure 2 provides a heatmap summarizing the accuracy of all models across the different datasets. It complements Table 1 in the full paper by visually reinforcing that convex models generally outperform their non-convex counterparts across nearly all datasets, with consistent and major improvements even for models like GATv2 and GraphGPS. This reaffirms the strength and generality of our convexification framework. For example, on MUTAG, the convex CGATv2 model reaches 87.5% accuracy, outperforming the two-layer (68.7%) and six-layer (70.0%) GINs. Similarly, CGraphGPS achieves 83.7%, exceeding the best non-convex variant by over 30%. This trend holds across other datasets like PTC-MR, PROTEINS, and NCI1, where convex versions of CGraphSAGE, CGATv2, and CGraphGPS outperform non-convex baselines by 10–40%.
I.2 Bar Plots
@@ I.3 Radar Plots
I.4 Experimental Results for Deeper Convex Models
Table 3 presents additional empirical results on Mutagenicity, comparing two-layer and six-layer convex versions of base models against their two-layer and six-layer non-convex counterparts [^1] As demonstrated in Section 5, two-layer convex models already exhibit superior performance compared to their two- and six-layer non-convex counterparts when all models incorporate DropOut, batch normalization, node input encoders, and edge feature encoders. To evaluate the performance of deeper GNN models in isolation, in this section we exclude these additional components.. Our results in Table 3 show that the accuracy of convex models can further improve with additional layers, which is consistent with our theoretical guarantees. Specifically, recall that Theorem 4.7 bounds the generalization error of Algorithm 2 for single-hidden-layer models. As specified in Section 4.5, our analysis extends layer-wise in the multi-layer setting, i.e., it applies to each hidden layer separately. The results in Table 3 therefore illustrate that errors remain controlled as depth increases. This offers strong empirical support for the practical effectiveness of multi-layer CGNNs, which is crucial since state-of-the-art GNN applications typically rely on deeper architectures.
Table 3: Accuracy ($\%\pm\sigma$) on Mutagenicity over 44 runs for two-layer and six-layer convex versions of base models against their two-layer and six-layer non-convex counterparts. Convex variants start with ’C’. A two-layer model is marked by (2L), while a six-layer one is marked by (6L). Results of convex models surpassing their non-convex variants by a large gap ($\sim$10–40%) are in bold, modest improvements are underlined, and improvements of six-layer convex models over the corresponding two-layer convex models are in dark green.
Model Mutagenicity
CGCN (2L) $54.7\pm 2.1$
CGCN (6L) $61.9\pm 2.7$
GCN (2L) $54.7\pm 1.1$
GCN (6L) $54.7\pm 1.9$
CGIN (2L) $56.8\pm 1.8$
CGIN (6L) $59.1\pm 1.8$
GIN (2L) $54.8\pm 2.0$
GIN (6L) $55.8\pm 2.0$
CGAT (2L) $55.7\pm 2.5$
CGAT (6L) $58.8\pm 3.6$
GAT (2L) $54.8\pm 2.0$
GAT (6L) $54.8\pm 2.0$
CGATv2 (2L) $56.6\pm 2.7$
CGATv2 (6L) $62.3\pm 4.2$
GATv2 (2L) $55.9\pm 1.4$
GATv2 (6L) $59.6\pm 6.2$
CResGatedGCN (2L) $56.2\pm 2.7$
CResGatedGCN (6L) $59.6\pm 3.7$
ResGatedGCN (2L) $56.1\pm 2.1$
ResGatedGCN (6L) $56.1\pm 2.1$
CGraphSAGE (2L) $55.1\pm 2.4$
CGraphSAGE (6L) $66.3\pm 5.3$
GraphSAGE (2L) $54.8\pm 2.0$
GraphSAGE (6L) $60.0\pm 3.3$
CGraphTrans (2L) $55.5\pm 2.3$
CGraphTrans (6L) $63.2\pm 4.7$
GraphTrans (2L) $54.8\pm 2.0$
GraphTrans (6L) $59.9\pm 4.6$
CGPS (2L) $57.7\pm 4.5$
CGPS (6L) $59.4\pm 4.1$
GPS (2L) $41.4\pm 8.7$
GPS (6L) $55.1\pm 5.5$
I.5 Experimental Results for DirGNNs
We herein provide additional empirical results for the recent baseline DirGNN (Rossi et al., 2024), which is a generic wrapper for computing graph convolution on directed graphs. If the graph is originally undirected, applying DirGNN effectively treats it as directed. In our simulations, we extend GCN, GAT and GraphSAGE with DirGNN, obtaining DirGCN, DirGAT and DirGraphSAGE, respectively. In Table 4, we report results on the MUTAG and PROTEINS datasets, where the results are generated as described in Table 1. Particularly, convex architectures are highlighted in bold. Our two-layer convex models consistently outperform their non-convex two- and six-layer counterparts by 10-40% accuracy.
Table 4: Accuracy ($\%\pm\sigma$) on MUTAG and PROTEINS over 44 runs for two-layer convex versions of DirGNNs against their two-layer and six-layer non-convex counterparts. Convex variants start with ’C’. A two-layer model is marked by (2L), while a six-layer one is marked by (6L). Results of convex models surpassing their non-convex variants by a large gap ($\sim$10–40%) are in bold, modest improvements are underlined, and improvements of six-layer convex models over the corresponding two-layer convex models are in dark green.
Model MUTAG PROTEINS
CDirGCN (2L) $\mathbf{85.0\pm 10.6}$ $\mathbf{72.5\pm 6.5}$
DirGCN (2L) $50.0\pm 10.0$ $45.1\pm 8.9$
DirGCN (6L) $57.5\pm 18.2$ $47.0\pm 10.6$
CDirGAT (2L) $\mathbf{75.0\pm 6.1}$ $\mathbf{73.6\pm 4.3}$
DirGAT (2L) $53.7\pm 16.7$ $59.7\pm 5.8$
DirGAT (6L) $61.2\pm 24.5$ $61.1\pm 10.5$
CDirGraphSAGE (2L) $\mathbf{81.2\pm 5.4}$ $\mathbf{74.3\pm 3.7}$
DirGraphSAGE (2L) $60.0\pm 14.5$ $39.2\pm 2.2$
DirGraphSAGE (6L) $41.2\pm 18.8$ $51.5\pm 9.4$
I.6 Why do Convexified Message-Passing GNNs Excel?
Next, we discuss why convexified GNNs perform well is valuable. We herein give a formal support for how each factor specified by the reviewer helps convexified models to excel, especially in small-data regimes.
  1. 1.
    Constraining filters by a nuclear-norm ball limits the hypothesis class and yields favorable generalization gap bounds. Specifically, under the conditions of Theorem 4.7, for each layer $\ell$ with graph filters $\mathcal{A}_{\ell}$ there exists a constant radius $B_{\ell}>0$ such that, if $\|\mathcal{A}_{\ell}\|_{\star}\leq B_{\ell}$, then the generalization gap scales as $\mathcal{O}(\frac{MC_{\sigma}(B_{\ell})}{\sqrt{m}})$ for a training set with $m$ samples and an $M$-Lipschitz continuous loss function, where $C_{\sigma}(B_{\ell})$ depends on the activation function $\sigma$. In contrast, unconstrained nonconvex training may effectively access much larger function classes (higher variance), which hurts generalization in small-data regimes. Our Theorem 4.7 thus indicates that nuclear-norm constraints control generalization for our CGNN class.
  2. 2.
    Nuclear-norm projection acts as spectral smoothing, removing or attenuating directions with small singular values. Projecting the learned filter onto a nuclear-norm ball effectively suppresses small singular-value directions of the filter matrix. In particular, constraining nuclear norms encourages solutions that spread the singular-value mass in a way that uses fewer very large singular values, as nuclear norm sums them up, and penalizes many moderate ones. This eventually encourages low-rank solutions, similarly to trace-norm regularization. On graph-structured data, small singular-value directions often correspond to high-frequency components in the graph spectrum. In graph signal processing terms, low-rank approximations correspond to low-dimensional subspaces which omit high-frequency noise (see, e.g., (Ramakrishna et al., 2020)). Thus, the nuclear-norm projection yields a smoother operator that suppresses noise and prevents overfitting to spurious high-frequency patterns. Namely, this explains the empirical gains we observe in low-data regimes.